# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ C Removal of highways (Type LG)
The behavior of Type L or LG models corroborates with empirical evidence for the local dispersion of economic agglomerations when interregional access improves. Baum-Snow-QJE2007 and Baum-Snow-et-al-REStat2017 presented evidence for US metro areas during 1950–1990 and Chinese prefectures during 1990–2010, respectively. These studies addressed the changes in the population or production size of the central area within the bigger region, both reporting a negative effect of improved interregional access. The local spreading of cities can also be interpreted as suburbanization in response to the improved intraurban transport infrastructure or diffusion of motorized transportation in classical Alonso–Muth–Mills theory. Structural change away from agriculture also frees up land around cities and contributes to urban density decline (Coeurdacier-etal-DP2022), which is also consistent with lowering the opportunity cost of land in the Alonso–Muth–Mills framework.
Future directions. We conclude by identifying several natural extensions of our framework that warrant further theoretical investigation.
@@
First, it is important to consider multiple types of mobile agents that differ in their proximity matrices and/or the degree of increasing returns they experience. Such heterogeneity is ubiquitous in multi-sector models (Fujita-et-al-EER1999; Hsu-EJ2012; Gaubert-AER2018; Davis-Dingel-JIE2020) and in intracity models with multiple types of agents (e.g., Fujita-Ogawa-RSUE1982; Lucas-Rossi-Hansberg-ECTA2002; Ahlfeldt-et-al-ECTA2015; Heblich-etal-QJE2020). Duranton-Morrow-Turner-REStud2014 studied the impact of new highway connections on intercity trade in the US and showed that heavier industries are more sensitive to improved access. The circular geography provides a tractable starting point for many-location analysis under such structures (Tabuchi-Thisse-JUE2011; Osawa-Akamatsu-DP2019). As Hsu-EJ2012 suggests, incorporating sectoral heterogeneity is also important for understanding the mechanisms behind the remarkable regularities in the size and spatial variation of cities (Mori-etal-PNAS2020).
@@
Second, models with a continuum of agents are complementary to granular spatial models (e.g., Ahlfeldt-et-al-CEPR2022), in which endogenous agglomeration arises from increasing returns and the indivisibility of agents. While continuum models can replicate systematic spatial regularities—such as periodic agglomeration patterns and city-size distributions including their fractal structure (e.g., Hsu-EJ2012; Tabuchi-Thisse-JUE2011; Mori-et-al-DP2022)—granular spatial models are better suited to capturing idiosyncratic location choices by superstar firms and large plants (e.g., Greenstone-et-al-JPE2010). Combining these two approaches may yield a deeper understanding of the spatial patterns of economic activities as outcomes of endogenous mechanisms.
@@ Appendix A Proofs
A.1   Proof of Lemma 1
Consider an infinitesimal migration shock at the uniform distribution $\bar{\bm{x}}$ so that the spatial distribution becomes $\bar{\bm{x}}+\bm{z}$. Let $T\equiv\{\bm{z}\in\mathbb{R}^{N}\mid\sum_{i\in\mathcal{I}}z_{i}=0\}$ be the set of all feasible deviations that preserves the total population.
The aggregate utility gain for migrants under a given deviation $\bm{z}$ can be measured by $\omega(\bm{z})\equiv\bm{z}^{\top}\mathbf{V}\bm{z}$, where $\top$ denotes transpose. To see this, let $\delta v_{i}(\bm{z})\equiv v_{i}(\bar{\bm{x}}+\bm{z})-v_{i}(\bar{\bm{x}})$ be the utility difference in each region under deviation $\bm{z}$. Then, the utility gain for a migrant from region $j$ to $i$ is $\frac{\bar{x}}{\bar{v}}\left(\delta v_{i}(\bm{z})-\delta v_{j}(\bm{z})\right)$. By adding up the utility gains for all the migrants in the economy, the aggregate utility gain for migrants is measured by $\omega(\bm{z})=\frac{\bar{x}}{\bar{v}}\sum_{i\in\mathcal{I}}\delta v_{i}(\bm{z})z_{i}=\frac{\bar{x}}{\bar{v}}\delta\bm{v}(\bm{z})^{\top}\bm{z}$, as there are $z_{i}$ agents migrated to region $i$ (if $z_{i}>0$) or from region $i$ (if $z_{i}<0$), each experiencing utility difference $\delta v_{i}(\bm{z})$ at both their origin and destination. The first-order approximation $\delta\bm{v}(\bm{z})\approx\bm{v}(\bar{\bm{x}})+\nabla\bm{v}(\bar{\bm{x}})\bm{z}-\bm{v}(\bar{\bm{x}})=\nabla\bm{v}(\bar{\bm{x}})\bm{z}$ gives $\omega(\bm{z})=\bm{z}^{\top}\mathbf{V}\bm{z}$.
The aggregate utility gain determines the stability of the symmetric equilibrium $\bar{\bm{x}}$. If $\omega(\bm{z})<0$ for all possible deviations, $\bar{\bm{x}}$ is stable; this is related to the concept of “regular ESS” (Sandholm-TE2010). If $\omega(\bm{z})>0$ for some deviation $\bm{z}$, $\bar{\bm{x}}$ is unstable. Thus, the maximum possible utility gain due to small migration shock is of interest.
[…]
A.2   Proof of Proposition 1
As discussed in the main text, we consider spatial models described by a payoff function (i.e., indirect utility) $\bm{v}(\bm{x})\equiv(v_{i}(\bm{x}))_{i\in\mathcal{I}}$, parametrized by a proximity matrix $[\phi_{ij}]$, along with Assumption C. Let $\mathbf{D}$ be the row-normalized proximity matrix, whose $(i,j)$th element is $\frac{\phi_{ij}}{\sum_{k\in\mathcal{I}}\phi_{ik}}$. Throughout, we assume that the payoff function $\bm{v}$ is differentiable if $x_{i}>0$ for all $i\in\mathcal{I}$, which holds true for all examples discussed in this paper. The precise version of the symmetry of exogenous local fundamentals in Assumption C is the following:
@@ A.2   Proof of Proposition 1
Suppose $N=4$. For instance, consider regions 1 and 3. Then, swapping their indices corresponds to applying the following permutation matrix:
@@
Under Assumption C, we can use the full dispersion as the initial state.
@@
[…]
@@
In Definition 1, a rational function $\Omega$ is a function of form $\Omega(\cdot)=\Omega^{\sharp}(\cdot)/\Omega^{\flat}(\cdot)$ with polynomials $\Omega^{\sharp}(\cdot)$ and $\Omega^{\flat}(\cdot)\neq 0$, where our convention is that $\Omega^{\flat}(\cdot)>0$. Given such $\Omega$, we let $\Omega(\mathbf{D})=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$, where, for a polynomial $P(\Theta)=c_{0}+c_{1}\Theta+c_{2}\Theta^{2}+\cdots$, we define $P(\mathbf{D})=c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}+\cdots$, with $\mathbf{I}$ being the identity matrix.
@@
[…]
@@
See Section A.3. ∎
We can assume, for example, the replicator dynamics (Taylor-Jorker-MB1978) to define local stability of spatial equilibria (see the proof of Lemma 4 for more examples).
Thus, we only need the eigenvalues of $\mathbf{V}$. A useful fact is that, if $\{(\Theta_{k},\bm{z}_{k})\}$ are the eigenpairs (eigenvalue–eigenvector pairs) of the normalized proximity matrix $\mathbf{D}$, then the eigenpairs of $\mathbf{V}=\Omega(\mathbf{D})$ are given by $\{(\Omega(\Theta_{k}),\bm{z}_{k})\}$ (e.g., Horn-Johnson-Book2012, Section 1.1). Thus, we need the eigenpairs of $\mathbf{D}$, and we have the following lemma.
@@
Assume Assumption C. The largest eigenvalue of $\mathbf{D}$ is $\Theta_{0}\equiv 1$, with $\bm{z}_{0}=(1,1,\ldots,1)$ being the associated eigenvector. Including $\Theta_{0}$, there are $\frac{N}{2}+1$ distinct eigenvalues. Every eigenvalue $\Theta_{k}$ ($k\neq 0$) is a strictly decreasing function of $\phi$, with $\lim_{\phi\downarrow 0}\Theta_{k}=1$ and $\lim_{\phi\uparrow 1}\Theta_{k}=0$. Let $\Theta_{\text{max}}$ denote the largest eigenvalue and $\Theta_{\text{min}}$ denote the smallest eigenvalue of $\mathbf{D}$ excluding $\Theta_{0}$, respectively. Then,
@@
As discussed in Section 2, if agglomeration (dispersion) force of the model is too strong, $\Omega(\Theta)>0$ ($\Omega(\Theta)<0$) can happen for all $\Theta$ whereby $\bar{\bm{x}}$ is unstable (stable) for all $\phi$. As we are interested in spatial agglomeration in the course of changing $\phi$, we assume that $\bar{\bm{x}}$ can switch its stability depending on $\phi$:
@@
Under Assumption E, we can define three prototypical classes of canonical models (see Fig. 4 in the main text for illustration).
@@
Under Assumption E, a canonical model with gain function $\Omega$ is
@@
    Type LG, if there can be two $\Theta^{*},\Theta^{**}\in(0,1)$ such that $\Omega(\Theta^{*})=\Omega(\Theta^{**})=0$ and $\Theta^{**}<\Theta^{*}$, with $\Omega(\Theta)<0$ for $\Theta\in(0,\Theta^{**})\cup(\Theta^{*},1)$ and $\Omega(\Theta)>0$ for Θ($\Theta\in(\Theta^{**},\Theta^{*})$.
@@
As we consider canonical models, there is a rational function $\Omega(\cdot)=\Omega^{\sharp}(\cdot)/\Omega^{\flat}(\cdot)$ with some polynomials $\Omega^{\sharp}$ and $\Omega^{\flat}(\cdot)>0$. That is, $\Omega^{\sharp}(\cdot)$ determines the sign of $\omega_{k}$ and thus governs the stability of $\bar{\bm{x}}$. We will focus on $\Omega^{\sharp}$ below, and let $\omega^{\sharp}\equiv\Omega^{\sharp}(\Theta_{k})$ so that $\operatorname{sgn}[\omega_{k}]=\operatorname{sgn}[\omega^{\sharp}_{k}]$. Fig. A.1 schematically shows connections between $\{\omega_{k}^{\sharp}\}$, $\Omega^{\sharp}(\Theta)$, and $\{\Theta_{k}\}$ to help understanding the following arguments.
Type L. By definition, there is $\Theta^{**}$ such that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(0,\Theta^{**})$, that $\Omega^{\sharp}(\Theta^{**})=0$, and that $\Omega^{\sharp}(\Theta^{**})>0$ for all $\Theta\in(\Theta^{**},1)$. Thus, $\bar{\bm{x}}$ is stable if and only if $\Theta_{k}\in(0,\Theta^{**})$, so that $\omega^{\sharp}_{k}=\Omega^{\sharp}(\Theta_{k})<0$, for all $k$, i.e., if $\Theta^{**}>\max_{k}\Theta_{k}=\Theta_{1}$. Thus, $\bar{\bm{x}}$ is stable for all $(\phi^{*},1)$ where ϕsuperscriptitalic-$\phi^{**}$ is the unique solution for $\Theta_{1}(\phi)=\Theta^{**}$. Because $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(\Theta^{**},1)$ and $\Theta_{1}$ is strictly decreasing, $\bar{\bm{x}}$ is unstable for all $(0,\phi^{**})$.
Type G. By definition, there is $\Theta^{*}$ such that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(\Theta^{*},1)$, that $\Omega^{\sharp}(\Theta^{*})=0$, and that $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$. By Lemma 5, $\{\Theta_{k}(\phi)\}$ are strictly decreasing from $1$. Thus, $\bar{\bm{x}}$ is stable if and only if $\Theta_{k}\in(\Theta^{*},1)$, so that $\omega^{\sharp}_{k}=\Omega^{\sharp}(\Theta_{k})<0$, for all $k$, i.e., if $\Theta^{*}<\min_{k}\Theta_{k}=\Theta_{N/2}$. Thus, $\bar{\bm{x}}$ is stable for all $(0,\phi^{*})$ where $\phi^{*}=(1-\sqrt{\Theta^{*}})/(1+\sqrt{\Theta^{*}})$ is the unique solution for $\Theta_{N/2}=\Theta^{*}$. Because $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$ and $\Theta_{N/2}$ is strictly decreasing, $\bar{\bm{x}}$ is unstable for all $(\phi^{*},1)$ because $\omega_{N/2}^{\sharp}>0$ for the range.
Type LG. Via similar logic, we see $\bar{\bm{x}}$ is stable if $\phi\in(0,\phi_{N/2}^{*})\cup(\phi_{1}^{*},1)$.
@@
Spatial patterns. Consider a state where $\bar{\bm{x}}$ is stable. From Lemma 4, at $\phi^{*}$ in Types G or LG, a polycentric pattern with $N/2$ peaks branches from $\bar{\bm{x}}$, while at ϕsuperscriptitalic-$\phi^{**}$ in Types L or LG, a monocentric configuration branches from $\bar{\bm{x}}$. $\square$
@@
The bifurcation towards the monocentric direction ($k=1$) is a double bifurcation, where the associated eigenvalue $\omega_{1}$ has multiplicity two—that is, two linearly independent eigenvectors exist. Migration patterns at this bifurcation take the form $c^{+}\bm{z}_{1}^{+}+c^{-}\bm{z}_{1}^{-}$ for $c^{+},c^{-}\in\mathbb{R}$. Under Assumption C, only $(c^{+},c^{-})=(c,0)$ or $(c,c)$ for some $c\in\mathbb{R}$ are admissible (ikeda2012spatial). While the choice between these two depends on the details of $\bm{v}$, this does not affect the conclusion of Proposition 1, as both combinations yield monocentric configurations. ∎
@@
Although Definition 2 introduces three prototypical model classes, Type LG models sometimes span multiple classes depending on parametric restrictions (e.g., models by Pfluger-Tabuchi-RSUE2010; Kucheryavyy-etal-JIE2024). In such cases, the parameter space can be partitioned to map model behavior to the typology. Also, the definition of models often impose parametric restrictions that fix their class. In principle, however, flexible specifications would allow empirical identification of the class supported by data through parameter estimation. This would pose a major methodological challenge, but also offers a promising direction for future research. ∎
@@
Assumption C assumes that $N$ is a multiple of four to ensure $\min_{k}\{\Theta_{k}\}=\Theta_{\frac{N}{2}}$. This is inconsequential for the broad implication of Proposition 1 on spatial patterns. If $N$ is an even, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{{\frac{N}{2}}-1},\Theta_{\frac{N}{2}}\}$. If $N$ is an odd, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{\lfloor\frac{N}{2}\rfloor},\Theta_{\lfloor\frac{N}{2}\rfloor-1}\}$. Thus, $\min_{k}\{\Theta_{k}\}$ corresponds to a polycentric direction, except for the case $N=2$ or $3$ in which polycentric patterns cannot occur. ∎
@@
First, no symmetry-breaking bifurcations can occur after the emergence of a single-peaked spatial pattern. For Type L models, the spatial configuration remains monocentric for the whole range of $\phi$ if the full dispersion is unstable (Figs. 8A, 1B and C.3)

### v8

@@ 5  Concluding remarks
In the intra-urban context, Baum-Snow-QJE2007 and Baum-Snow-et-al-REStat2017 provide evidence for the US metropolitan areas from 1950 to 1990 and for Chinese prefectures from 1990 to 2010. Both studies examine how the share of population or production in the central area within a larger region changes as the transportation infrastructure expands, and both report negative effects in the central area. This is consistent with the behavior of Type L or LG models after transport investments. Such local spread can also be viewed as suburbanization driven by improved intra-urban transport infrastructure or by the diffusion of motorized transportation in the Alonso–Muth–Mills framework. [^24] The structural transformation away from agriculture frees land around cities and also contributes to the decline of urban density (Coeurdacier-etal-DP2022). This can also be interpreted through the AMM framework as a reduction in the opportunity cost of land.
As seen above, a unified theoretical framework can help synthesize and interpret empirical findings. Since this study focuses on static models with a single agent type, further scrutinies are essential for providing a bird’s-eye view of both the empirical evidence and the now vast quantitative spatial economics literature. We conclude by outlining two directions that merit further theoretical investigation.
@@
First, it is important to consider multiple types of mobile agents that differ in their proximity matrices and/or the degree of increasing returns they experience. Such heterogeneity is ubiquitous in multi-sector models (Fujita-et-al-EER1999; Hsu-EJ2012; Gaubert-AER2018; Davis-Dingel-JIE2020) and in intracity models with multiple types of agents (e.g., Fujita-Ogawa-RSUE1982; Lucas-Rossi-Hansberg-ECTA2002; Ahlfeldt-et-al-ECTA2015; Heblich-etal-QJE2020). For example, Duranton-Morrow-Turner-REStud2014 studied the impact of new highway connections on intercity trade in the US and showed that heavier industries are more sensitive to improved access. Allen-etal-AER2024 considered a quite general spatial model with multiple spatial interactions, but the characterization of endogenous equilibrium spatial structure has yet to be done, in particular for the cases with multiple equilibria. Circular geography provides a tractable starting point for analysis of many-locations under such structures (Tabuchi-Thisse-JUE2011; Osawa-Akamatsu-DP2019). As Hsu-EJ2012 suggests, the incorporation of sectoral heterogeneity can be particularly important for understanding the mechanisms behind the remarkable regularities in the size and spatial variation of cities (Mori-etal-PNAS2020).
@@
Second, models with a continuum of agents as considered in this study are complementary to “granular” spatial models (e.g., Ahlfeldt-et-al-CEPR2022), in which endogenous agglomeration arises from increasing returns and the indivisibility of agents. Continuum-agent models can replicate systematic spatial regularities, such as periodic agglomeration patterns and city-size distributions that include their fractal structure (e.g., Hsu-EJ2012; Tabuchi-Thisse-JUE2011; Mori-et-al-DP2022). Granular spatial models are better suited to capture idiosyncratic location choices by superstar firms and large plants (e.g., Greenstone-et-al-JPE2010). Combining these two approaches may yield a deeper understanding of the spatial patterns of economic activities as the result of endogenous forces.
@@ Appendix A Proofs
A.1  Proof of Lemma˜1
First, we can represent $\Delta$ in terms of marginal migration from region $2$ to $1$: $\Delta(\epsilon)\coloneq\Delta(\bar{\bm{x}}+\frac{\epsilon}{2}(1,-1))=v_{1}(\bar{x}+\frac{\epsilon}{2},\bar{x}-\frac{\epsilon}{2})-v_{2}(\bar{x}+\frac{\epsilon}{2},\bar{x}-\frac{\epsilon}{2})$. Then, $\omega=\frac{\bar{x}}{\bar{v}}\Delta^{\prime}(\epsilon)|_{\epsilon=0}$, meaning that $\omega$ is proportional to the directional derivative of $\Delta$ with respect to the unit migration from region $2$ to $1$, as $\frac{1}{2}(1,-1)$ is a normalized vector. Here, we used the fact that $\frac{\partial{v_{1}}}{\partial{x_{1}}}(\bar{\bm{x}})=\frac{\partial{v_{2}}}{\partial{x_{2}}}(\bar{\bm{x}})$ and $\frac{\partial{v_{1}}}{\partial{x_{2}}}(\bar{\bm{x}})=\frac{\partial{v_{2}}}{\partial{x_{1}}}(\bar{\bm{x}})$ due to the symmetry of the regions at $\bar{\bm{x}}$. The stability condition based on the sign of $\omega$ is valid for general $\bm{v}$ provided that $\bar{v}>0$.
A.2  Proof of Lemma˜2
Let $\delta v_{i}(\bm{z})\equiv v_{i}(\bar{\bm{x}}+\bm{z})-v_{i}(\bar{\bm{x}})$ be the utility difference in each region under deviation $\bm{z}$. The utility gain for a mover from region $j$ to $i$ is
$\frac{\bar{x}}{\bar{v}}\left(\delta v_{i}(\bm{z})-\delta v_{j}(\bm{z})\right).$ (A.1)
By adding up the utility gains for all the migrants in the economy, the aggregate utility gain for migrants is measured by $\omega(\bm{z})=\frac{\bar{x}}{\bar{v}}\sum_{i\in\mathcal{I}}\delta v_{i}(\bm{z})z_{i}=\frac{\bar{x}}{\bar{v}}\delta\bm{v}(\bm{z})^{\top}\bm{z}$, as there are $z_{i}$ agents migrated to region $i$ (if $z_{i}>0$) or from region $i$ (if $z_{i}<0$), each experiencing utility difference $\delta v_{i}(\bm{z})$ at both their origin and destination. The first-order approximation shows $\delta\bm{v}(\bm{z})\approx\bm{v}(\bar{\bm{x}})+\nabla\bm{v}(\bar{\bm{x}})\bm{z}-\bm{v}(\bar{\bm{x}})=\nabla\bm{v}(\bar{\bm{x}})\bm{z}$ and hence gives $\omega(\bm{z})=\bm{z}^{\top}\mathbf{V}\bm{z}$.
It is noted that $\omega(\bm{z})$ must be considered subject to $\bm{z}\in T$, where $T\equiv\{\bm{z}\in\mathbb{R}^{N}\mid\sum_{i\in\mathcal{I}}z_{i}=0\}$ is the set of all feasible deviations that preserves the total population. To avoid technicality, the main text do not mention the constraint $\bm{z}_{k}\in T$. This constraint excludes deviations of the form $\bm{z}=(\epsilon,\epsilon,\ldots,\epsilon)$, which represents the symmetric increase or decrease of population across all regions.
A.3  Proof of Proposition˜1
We consider spatial models described by a payoff function (i.e., indirect utility) $\bm{v}(\bm{x})\equiv(v_{i}(\bm{x}))_{i\in\mathcal{I}}$, parametrized by a proximity matrix $\phi_{ij}]$, along with Assumption˜C. Let $\mathbf{D}$ be the row-normalized proximity matrix, whose $(i,j)$th element is $\frac{\phi_{ij}}{\sum_{k\in\mathcal{I}}\phi_{ik}}$. Throughout, we assume that $\bm{v}$ is differentiable if $x_{i}>0$ for all $i\in\mathcal{I}$. The precise version of the symmetry of exogenous local fundamentals in Assumption˜C is the following:
@@ A.3  Proof of Proposition˜1
Suppose $N=4$. If we consider regions 1 and 3, swapping their indices corresponds to applying the following permutation matrix to the spatial distribution and the payoff function:
@@
Under Assumption˜C, we can use the full dispersion as the initial state.
@@
Under Assumption˜C (including Assumption˜S), the uniform distribution of agents, $\bar{\bm{x}}=(\bar{x},\bar{x},\ldots,\bar{x})$ with $\bar{x}\equiv 1/N$, is a spatial equilibrium. ∎
@@
In Definition˜2, a rational function $\Omega$ is a function of form $\Omega(\cdot)=\Omega^{\sharp}(\cdot)/\Omega^{\flat}(\cdot)$ with polynomials $\Omega^{\sharp}(\cdot)$ and $\Omega^{\flat}(\cdot)\neq 0$, where our convention is that $\Omega^{\flat}(\cdot)>0$. Given such $\Omega$, we let $\Omega(\mathbf{D})=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$, where, for a polynomial $P(\Theta)=c_{0}+c_{1}\Theta+c_{2}\Theta^{2}+\cdots$, we define $P(\mathbf{D})=c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}+\cdots$, with $\mathbf{I}$ being the identity matrix.
@@
Assume a canonical model and assume Assumption˜C. For a wide class of myopic adjustment dynamics, $\bar{\bm{x}}$ is stable (unstable) if the largest eigenvalue of $\mathbf{V}$, excluding the one corresponding to $\bm{1}=(1,1,\ldots,1)$, is smaller (greater) than zero. Furthermore, if only the sign of the largest eigenvalue turns from negative to positive at some $\phi^{*}\in(0,1)$, then a new spatial equilibrium branches from $\bar{\bm{x}}$ at $\phi^{*}$, toward the direction of associated eigenvector. ∎
@@
See Section˜A.4. ∎
We can assume, for example, the replicator dynamics (Taylor-Jorker-MB1978) to define local stability of spatial equilibria (see the proof of Lemma˜4 for more examples).
Thus, we only need the eigenvalues of $\mathbf{V}$. A useful fact is that, if $\{(\Theta_{k},\bm{z}_{k})\}$ are the eigenpairs (eigenvalue–eigenvector pairs) of the normalized proximity matrix $\mathbf{D}$, then the eigenpairs of $\mathbf{V}=\Omega(\mathbf{D})$ are given by $\{(\Omega(\Theta_{k}),\bm{z}_{k})\}$ (e.g., Horn-Johnson-Book2012, Section 1.1). Thus, we need the eigenpairs of $\mathbf{D}$. For the sake of simplicity, we assume that $N$ is a multiple of four. Then, we have the following lemma.
@@
Assume Assumption˜C. The largest eigenvalue of $\mathbf{D}$ is $\Theta_{0}\equiv 1$, with $\bm{z}_{0}=(1,1,\ldots,1)$ being the associated eigenvector. Including $\Theta_{0}$, there are $\frac{N}{2}+1$ distinct eigenvalues. Every eigenvalue $\Theta_{k}$ ($k\neq 0$) is a strictly decreasing function of $\phi$, with $\lim_{\phi\downarrow 0}\Theta_{k}=1$ and $\lim_{\phi\uparrow 1}\Theta_{k}=0$. Let $\Theta_{\text{max}}$ denote the largest eigenvalue and $\Theta_{\text{min}}$ denote the smallest eigenvalue of $\mathbf{D}$ excluding $\Theta_{0}$, respectively. Further, assume that $N$ is a multiple of four. Then,
@@
As discussed in Section˜2, if agglomeration (dispersion) force of the model is too strong, $\Omega(\Theta)>0$ ($\Omega(\Theta)<0$) can happen for all $\Theta$ whereby $\bar{\bm{x}}$ is unstable (stable) for all $\phi$. As we are interested in spatial agglomeration in the course of changing $\phi$, we assume that $\bar{\bm{x}}$ can switch its stability depending on $\phi$:
@@
Under Assumption˜E, we can define three prototypical classes of canonical models (see Fig.˜5 in the main text for illustration).
@@
Under Assumption˜E, a canonical model with gain function $\Omega$ is
@@
    Type LG, if there can be two $\Theta^{*},\Theta^{**}\in(0,1)$ such that $\Omega(\Theta^{*})=\Omega(\Theta^{**})=0$ and $\Theta^{**}<\Theta^{*}$, with $\Omega(\Theta)<0$ for $\Theta\in(0,\Theta^{**})\cup(\Theta^{*},1)$ and $\Omega(\Theta)>0$ for $\Theta\in(\Theta^{**},\Theta^{*})$.
@@
As we consider canonical models, there is a rational function $\Omega(\cdot)=\Omega^{\sharp}(\cdot)/\Omega^{\flat}(\cdot)$ with some polynomials $\Omega^{\sharp}$ and $\Omega^{\flat}(\cdot)>0$. That is, $\Omega^{\sharp}(\cdot)$ determines the sign of $\omega_{k}$ and thus governs the stability of $\bar{\bm{x}}$. We will focus on $\Omega^{\sharp}$ below, and let $\omega^{\sharp}\equiv\Omega^{\sharp}(\Theta_{k})$ so that $\operatorname{sgn}[\omega_{k}]=\operatorname{sgn}[\omega^{\sharp}_{k}]$. Fig.˜A.1 schematically shows connections between $\{\omega_{k}^{\sharp}\}$, $\Omega^{\sharp}(\Theta)$, and $\{\Theta_{k}\}$ to help understanding the following arguments.
Type L.
By definition, there is $\Theta^{**}$ such that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(0,\Theta^{**})$, that $\Omega^{\sharp}(\Theta^{**})=0$, and that $\Omega^{\sharp}(\Theta^{**})>0$ for all $\Theta\in(\Theta^{**},1)$. Thus, $\bar{\bm{x}}$ is stable if and only if $\Theta_{k}\in(0,\Theta^{**})$, so that $\omega^{\sharp}_{k}=\Omega^{\sharp}(\Theta_{k})<0$, for all $k$, i.e., if $\Theta^{**}>\max_{k}\Theta_{k}=\Theta_{1}$. Thus, $\bar{\bm{x}}$ is stable for all $(\phi^{*},1)$ where $\phi^{**}$ is the unique solution for $\Theta_{1}(\phi)=\Theta^{**}$. Because $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(\Theta^{**},1)$ and $\Theta_{1}$ is strictly decreasing, $\bar{\bm{x}}$ is unstable for all $(0,\phi^{**})$.
Type G.
By definition, there is $\Theta^{*}$ such that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(\Theta^{*},1)$, that $\Omega^{\sharp}(\Theta^{*})=0$, and that $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$. By Lemma˜5, $\{\Theta_{k}(\phi)\}$ are strictly decreasing from $1$. Thus, $\bar{\bm{x}}$ is stable if and only if $\Theta_{k}\in(\Theta^{*},1)$, so that $\omega^{\sharp}_{k}=\Omega^{\sharp}(\Theta_{k})<0$, for all $k$, i.e., if $\Theta^{*}<\min_{k}\Theta_{k}=\Theta_{N/2}$. Thus, $\bar{\bm{x}}$ is stable for all $(0,\phi^{*})$ where $\phi^{*}=(1-\sqrt{\Theta^{*}})/(1+\sqrt{\Theta^{*}})$ is the unique solution for $\Theta_{N/2}=\Theta^{*}$. Because $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$ and $\Theta_{N/2}$ is strictly decreasing, $\bar{\bm{x}}$ is unstable for all $(\phi^{*},1)$ because $\omega_{N/2}^{\sharp}>0$ for the range.
Type LG.
Via similar logic, we see $\bar{\bm{x}}$ is stable if $\phi\in(0,\phi_{N/2}^{*})\cup(\phi_{1}^{*},1)$.
@@
Spatial patterns.
Consider a state where $\bar{\bm{x}}$ is stable. From Lemma˜4, at $\phi^{*}$ in Types G or LG, a polycentric pattern with $N/2$ peaks branches from $\bar{\bm{x}}$, while at $\phi^{**}$ in Types L or LG, a monocentric configuration branches from $\bar{\bm{x}}$. $\square$
@@
The bifurcation towards the monocentric direction ($k=1$) is a double bifurcation, where the associated eigenvalue $\omega_{1}$ has multiplicity two, that is, there are two linearly independent eigenvectors. Migration patterns at this bifurcation take the form $c^{+}\bm{z}_{1}^{+}+c^{-}\bm{z}_{1}^{-}$ for $c^{+},c^{-}\in\mathbb{R}$. Under Assumption˜C, only $(c^{+},c^{-})=(c,0)$ or $(c,c)$ for some $c\in\mathbb{R}$ are admissible (ikeda2012spatial). The conclusion of Proposition˜1 is not affected as both combinations yield monocentric configurations. ∎
@@
Although Definition˜3 introduces three prototypical model classes, Type LG models sometimes span multiple classes depending on parametric restrictions (e.g., models by Pfluger-Tabuchi-RSUE2010; Kucheryavyy-etal-JIE2024). In such cases, the parameter space can be partitioned to map model behavior to the typology. Also, the definition of models often impose parametric restrictions that fix their class. In principle, flexible specifications would allow empirical identification of the class supported by data through parameter estimation. ∎
@@
In Lemma˜5, we assume that $N$ is a multiple of four to ensure $\min_{k}\{\Theta_{k}\}=\Theta_{\frac{N}{2}}$. This is inconsequential for the broad implication of Proposition˜1 on spatial patterns. If $N$ is an even, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{{\frac{N}{2}}-1},\Theta_{\frac{N}{2}}\}$. If $N$ is an odd, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{\lfloor\frac{N}{2}\rfloor},\Theta_{\lfloor\frac{N}{2}\rfloor-1}\}$. Thus, $\min_{k}\{\Theta_{k}\}$ corresponds to a polycentric direction, except for the case $N=2$ or $3$ in which polycentric patterns cannot occur. ∎
@@
First, no symmetry-breaking bifurcations can occur after the emergence of a single-peaked spatial pattern. For Type L models, the spatial configuration remains monocentric for the whole range of $\phi$ if the full dispersion is unstable (Figs.˜9A, 1B and C.3)